# Startup Acquisitions: Acquihires and Talent Hoarding

### v3

@@ 4 Hiring, Separation, and Unemployment
The increase in hiring follows immediately because in the presence of talent hoarding, not only high-match firms but also low-match firms may pursue an acquihire. The increase in separation and unemployment is more subtle. Essentially, when either the correlation between firms’ adverse shocks or or the (marginal) probability of suffering a downgrade $\gamma$ is sufficiently high, talent hoarding raises separation and unemployment. In the case in which $r$ is high, this is because firm 1’s shock is informative of firm 2’s shock because of the correlation, hence allowing firm 1, whenever it draws a negative shock, to forgo the costly talent hoarding in the second period. Similarly, when $\gamma$ is sufficiently high, firm 1 can be fairly confident that whenever a downturn occurs, the competitor will be downgraded, once more prompting the firm to forgo talent hoarding. Further, when $\gamma$ or $r$ is sufficiently high, firm 1 is often right in laying off the entrepreneur, as firm 2 will indeed have a low match, which in turn will lead to unemployment. Finally, all statements in the proposition are strict whenever $\lambda$ is sufficiently high so that any talent hoarding at all takes place.
@@ 5 Discussion and Conclusion
We have presented a simple, yet general, reduced-form model of startup acquisitions. We showed that acquihires may not only reflect firms’ desire to hire talented employees but also be rooted in an incentive to engage in inefficient talent hoarding, thus potentially warranting regulators’ attention. Further, we showed that acquihires may decrease consumer surplus and increase job volatility of acquihired employees, thereby giving further reasons for regulatory scrutiny.
Our baseline model relies on several simplifying assumptions. As we discuss below, the main results hold even if we relax some of those assumptions. Moreover, the extensions of our baseline model reveal several additional insights. The formal analysis and results can be found in the Online Appendix.
People and technology.
In our baseline model, the startup’s value lies solely in its employees. However, some startups also possess valuable technology. We posit that the key distinction between people and technology is that firms can sell or license acquired technology but not employees. [^13] For a discussion of technology markets, see, e.g., Arora et al. (2001) and Gans and Stern (2000). We extend the model so that some share of the startup’s value is also due to its technology. An acquirer can trade the startup’s technology (but not the employees) to its competitor. As we show, a low-match acquirer indeed has an incentive to sell the technology to a high-match competitor because the price compensates her for any decrease in profit due to a more efficient competitor. Thus firms do not have an incentive to hoard technology, while the incentive to hoard talent remains. Interestingly, talent hoarding now occurs for a strictly larger set of parameters, as the option to resell technology effectively subsidizes talent hoarding.
@@
Partial acquisitions.
Instead of integrating the startup after acquiring it, the acquirer could allow it to continue operating independently in its own market. Moreover, instead of buying the startup outright, a firm could acquire a partial stake in it. In this context, the key questions are how the startup’s profits and control rights are allocated. First, we assume that following an investment the entrepreneur receives a dividend and a wage, while an investor only receives a dividend as their share of the startup’s profits. To microfound these payoffs, we assume that the presence of an outside investor gives rise to an agency problem, which we capture in reduced form, thus allowing for a wide range of agency models while maintaining a simple and (relatively) tractable setup. Second, considering control rights, we are primarily interested in the investor’s ability to block and the entrepreneur’s ability to push through an acquihire by the investor’s competitor. We assume that the entrepreneur can always block an acquihire, as she could sell her shares but refuse to work for the acquiring firm. If the entrepreneur would like to be acquihired and the investor does not want the acquihire to go through, the entrepreneur can (in the spirit of typical shareholder agreements) try to “drag along” the investor and force the transaction. We assume that the investor has a probabilistic chance of blocking this attempt and merely impose that the probability of successfully blocking an acquihire is increasing in the investor’s stake in the startup.
We find that an acquihire can be more profitable than buying a startup and letting it operate independently, implying that talent hoarding persists in this extension. With sufficiently strong blocking rights, an investment may constitute a viable and cheaper alternative to an acquihire. Therefore, as investments are cheaper than acquihires, the possibility of partial ownership may reduce the frequency of acquihires while increasing the frequency with which some transaction takes place. Notably, the reduction in overall profits is lower in the case of investments than the case of low-match acquihires. Therefore, the possibility of investments is more likely to lead to inefficient market outcomes, albeit at a lower degree of inefficiency.
@@
Appendix A Proof of Proposition 3
@@
Absent incentives to hoard talent, only high-match firms do an acquihire. Thus, a layoff only takes place if the economy enters a downturn and the period-1 acquirer is hit by an adverse shock. Further, following a layoff, we observe unemployment only if the competitor was a low-match firm or if it was a high-match firm that got hit by an adverse shock. Finally, hiring takes place in period 1 unless both firms are low-match. Taken together, the probability of observing a layoff in period 2 is
@@
and the probability of observing unemployment in period 2 is
@@
First, note that period-2 incentives coincide with those in period 1 absent an economic downturn, ruling out layoffs and unemployment. Following a downturn, three cases arise in period 2:
@@
To compare hiring, separation, and unemployment, we need to consider three cases.
@@
Case 1: $\lambda(1-r)(1-\gamma)>\lambda_{A}$. Then, $\lambda\geq\lambda_{A}^{1}=\lambda_{A}\frac{2}{2-\gamma\delta}$. Hence, a low-match firm will do an acquihire in both periods so no layoffs or unemployment are observed, which is less than in the benchmark. Thus, irrespective of whether the economy hits a downturn, the entrepreneur is always employed without separation.
@@
Case 2: $\lambda(1-\gamma(1-r))>\lambda_{A}>\lambda(1-r)(1-\gamma)$. Hence, $\lambda>\lambda_{A}^{2}$ so firm 1 will always do an acquihire in period 1. In period 2, firm 1 will maintain employment of the entrepreneur unless it receives shock $D$. Hence, the probability of a layoff will be $\delta\gamma$, which is larger than the benchmark layoff rate $l^{*}$. Moreover, the probability of transition to unemployment will be $\delta(\gamma-\lambda P(D,N)),$ which is larger than $u^{*}$ iff $\frac{1-\lambda}{\lambda}>(1-r)(1-\gamma)$.
@@
Case 3: $\lambda_{A}>\lambda(1-\gamma(1-r))$. If $\lambda>\lambda_{A}^{3}$ then firm 1 will do an acquihire in period 1. Firm 1 will maintain employment in period 2 unless it was hit by a downturn and has a low match. Hence, the probability of observing a layoff is $\delta(\gamma+(1-\lambda)(1-\gamma))=\delta(1-\lambda(1-\gamma))$, which is larger than $l^{*}$. The probability of observing a transition to unemployment is $\delta([2\lambda-\lambda^{2}]\gamma-\lambda^{2}P(D,N)+(1-\lambda)^{2})$ which is larger than $u^{*}$. If instead $\lambda<\lambda_{A}^{3}$, then we have no talent hoarding in either stage and the equilibrium is identical to the benchmark.
@@
Finally, observe that the condition $\min\left\{\frac{\lambda_{A}}{\lambda},\frac{1-\lambda}{\lambda}\right\}>(1-r)(1-\gamma)$ implies that we are either in Case 2 or 3. Further, in Case 2, it ensures that we have more unemployment than in the benchmark.
Appendix B Online Appendix: Not for Publication
The online appendix covers the extensions discussed in the conclusion of the paper, as well as the robustness exercises concerning the timing and surplus sharing in the baseline model mentioned in footnote 6.
B.1 People and Technology
Consider a situation where the total value of the startup consists of the people who work for the startup and the technology owned by the startup. The fundamental difference between the employees and the technology, from the acquirer’s point of view, is that technology can be sold (or licensed), while the people cannot. In our model, this implies that the acquirer can resell the startup’s technology to the competitor, whenever such a sale increases joint profits. Suppose that the share of the value of the startup generated by the technology is $\delta\in[0,1]$, while the share generated by employees is $(1-\delta)$. Moreover, for simplicity, assume that acquiring just the technology (or just the employees) generates $\delta$ (or $1-\delta$) of the impact that acquiring the entire startup would have. Just as before, a firm can have a high match value with the startup with probability $\lambda$, where we assume for simplicity that the match value of the startup to a firm applies to both the people and the technology identically. The match value of the firm is private information at the beginning of the game. The timing of the game in stage 1 is:
  1. 1.
  2. 2.
    The startup accepts or rejects the bid.
  3. 3.
Stage 2 is like stage 1 but the roles of firms 1 and 2 are reversed. To accommodate the possibility of selling the startup’s technology, we slightly adapt the notation from the main text. Suppose firm 1 with match $\theta_{1}$ did an acquisition at price $p$. Absent any sale of the startup’s technology, profits read
$\text{Firm}~{}1:$ $\;\Pi_{F}+\bar{\pi}_{F}^{\theta_{1}}-p$
$\text{Firm}~{}2:$ $\;\Pi_{F}-\underline{\pi}_{F}^{\theta_{1}}$
[…]
$\text{Firm}~{}1:$ $\;\Pi_{F}+(1-\delta)\bar{\pi}_{F}^{\theta_{1}}-\delta\underline{\pi}_{F}^{\theta_{2}}-p+q$
$\text{Firm}~{}2:$ $\;\Pi_{F}-(1-\delta)\underline{\pi}_{F}^{\theta_{1}}+\delta\bar{\pi}_{F}^{\theta_{2}}-q$
[…]
$\Pi_{F}$ $+(1-\delta)\bar{\pi}_{F}^{\theta_{1}}-\delta\underline{\pi}_{F}^{\theta_{2}}-p+\Pi_{F}-(1-\delta)\underline{\pi}_{F}^{\theta_{1}}+\delta\bar{\pi}_{F}^{\theta_{2}}-\left(\Pi_{F}+\bar{\pi}_{F}^{\theta_{1}}-p+\Pi_{F}-\underline{\pi}_{F}^{\theta_{1}}\right)$
$=$ $\delta\left(\bar{\pi}_{F}^{\theta_{2}}+\underline{\pi}_{F}^{\theta_{1}}-\bar{\pi}_{F}^{\theta_{1}}-\underline{\pi}_{F}^{\theta_{2}}\right).$
The case we are interested in, is when a low-match firm 1 sells technology to a high-match firm 2. Then, the surplus reads $\delta\left(\bar{\pi}_{F}^{H}+\underline{\pi}_{F}^{L}-\bar{\pi}_{F}^{L}-\underline{\pi}_{F}^{H}\right)$. We now make two assumptions on profits. The first corresponds to Assumption 1 in the baseline model in the main text (in this extension’s notation). The second ensures that the surplus resulting from a technology sale from a low-match to a high-match firm is positive.
Assumption B1.
[…]
Assumption B2.
Proposition B1.
Under Assumptions B1 and B2, firm 1’s behavior in any PBE is uniquely specified. Namely, if firm 1 has a high match, it will make an acquisition and not sell the technology; if it has a low match, it will make an acquisition and sell the startup’s technology to a high-match (but not a low-match) firm 2 if and only if
$\lambda\geq\lambda_{A}(\delta)\equiv\frac{\pi_{E}-\bar{\pi}_{F}^{L}}{\underline{\pi}_{F}^{H}+\frac{\delta}{2}\left(\bar{\pi}_{F}^{H}+\underline{\pi}_{F}^{L}-\bar{\pi}_{F}^{L}-\underline{\pi}_{F}^{H}\right)}$
and do nothing otherwise.
Proof.

### v4

@@ 4 Hiring and Layoffs
The increase in hiring follows immediately because in the presence of talent hoarding, not only high-match firms but also low-match firms may pursue an acquihire. The increase in layoffs is more subtle. Essentially, when either the correlation between firms’ adverse shocks or the (marginal) probability of suffering a downgrade $\gamma$ is sufficiently high, talent hoarding raises layoffs. In the case in which $r$ is high, this is because firm 1’s shock is informative of firm 2’s shock because of the correlation, hence allowing firm 1, whenever it draws a negative shock, to forgo the costly talent hoarding in the second period. Similarly, when $\gamma$ is sufficiently high, firm 1 can be fairly confident that whenever a downturn occurs, the competitor will be downgraded, once more prompting the firm to forgo talent hoarding. Further, when $\gamma$ or $r$ is sufficiently high, firm 1 is often right in laying off the entrepreneur, as firm 2 will indeed have a low match, which in turn will lead to the entrepreneur exiting the industry. Finally, all statements in the proposition are strict whenever $\lambda$ is sufficiently high so that any talent hoarding at all takes place.
After Google announced they would be laying off 12,000 people in 2023 (and following layoffs at other tech companies), a tech commentator John Gruber wrote the following: “There are numerous reasons the tech industry wound up at this layoffpalooza, but I think the main reason is that the biggest companies got caught up in a game where they tried to hire everyone, whether they needed them or not, to keep talent away from competitors and keep talent away from small upstarts (or from founding their own small upstarts). These big companies were just hiring to hire, and now the jig is up.” [^14] From https://daringfireball.net/linked/2023/01/20/google-layoffs, accessed on June 28, 2024. Sam Lessin, the founder of Drop.io mentioned in the introduction, expressed a similar sentiment in an interview with CNBC. [^15] Available at https://www.cnbc.com/video/2024/02/06/tech-companies-were-hoarding-the-talent-says-slow-ventures-sam-lessin.html, accessed on June 28, 2024. Thus, the main economic forces that our model identifies align with industry insiders’ observations.
5 Other Extensions
We now extend our baseline model in two directions, both of which lead to more talent hoarding. First, we allow for the situation in which the startup is valuable to the firms not only because of its employees but also because of its technology. In doing so, we emphasize the difference between labor and other inputs in that firms do not acquire property rights over labor. Thus, in contrast to the labor, the technology can be sold (or licensed) after an acquisition. Second, we consider partial acquisitions, or investments, into startups. In this case, the acquirer does not acquire the startup in its entirety but only a fraction of it and in turn receives a fraction of the startup’s profits as well as some decision power. We show that when the decision power suffices to block acquisitions by competitors, such investments constitute an attractive alternative to full acquihires. As it turns out, the possibility of investments is more likely to lead to inefficient market outcomes, albeit at a lower degree of inefficiency.
5.1 People and Technology
Consider a situation where the total value of the startup consists of the people who work for the startup and the technology owned by the startup. The fundamental difference between the employees and the technology, from the acquirer’s point of view, is that technology can be sold (or licensed), while the people cannot. In our model, this implies that the acquirer can resell the startup’s technology to the competitor, whenever such a sale increases joint profits. Suppose that the share of the value of the startup generated by the technology is $\tau\in[0,1]$, while the share generated by employees is $(1-\tau)$. Moreover, for simplicity, assume that acquiring just the technology (or just the employees) generates $\tau$ (or $1-\tau$) of the impact that acquiring the entire startup would have. Just as before, a firm can have a high match value with the startup with probability $\lambda$, where we assume for simplicity that the match value of the startup to a firm applies to both the people and the technology identically. The match value of the firm is private information at the beginning of the game. The timing of the game in stage 1 is:
  1. 1.
  2. 2.
    The startup accepts or rejects the bid.
  3. 3.
Stage 2 is like stage 1 but the roles of firms 1 and 2 are reversed. To accommodate the possibility of selling the startup’s technology, we slightly adapt the notation from the baseline model. Suppose firm 1 with match $\theta_{1}$ did an acquisition at price $p$. Absent any sale of the startup’s technology, profits read
$\text{Firm}~{}1:$ $\;\Pi_{F}+\bar{\pi}_{F}^{\theta_{1}}-p$
$\text{Firm}~{}2:$ $\;\Pi_{F}-\underline{\pi}_{F}^{\theta_{1}}$
[…]
$\text{Firm}~{}1:$ $\;\Pi_{F}+(1-\tau)\bar{\pi}_{F}^{\theta_{1}}-\tau\underline{\pi}_{F}^{\theta_{2}}-p+q$
$\text{Firm}~{}2:$ $\;\Pi_{F}-(1-\tau)\underline{\pi}_{F}^{\theta_{1}}+\tau\bar{\pi}_{F}^{\theta_{2}}-q$
[…]
$\Pi_{F}$ $+(1-\tau)\bar{\pi}_{F}^{\theta_{1}}-\tau\underline{\pi}_{F}^{\theta_{2}}-p+\Pi_{F}-(1-\tau)\underline{\pi}_{F}^{\theta_{1}}+\tau\bar{\pi}_{F}^{\theta_{2}}-\left(\Pi_{F}+\bar{\pi}_{F}^{\theta_{1}}-p+\Pi_{F}-\underline{\pi}_{F}^{\theta_{1}}\right)$
$=$ $\tau\left(\bar{\pi}_{F}^{\theta_{2}}+\underline{\pi}_{F}^{\theta_{1}}-\bar{\pi}_{F}^{\theta_{1}}-\underline{\pi}_{F}^{\theta_{2}}\right).$
The case we are interested in, is when a low-match firm 1 sells technology to a high-match firm 2. Then, the surplus reads $\tau\left(\bar{\pi}_{F}^{H}+\underline{\pi}_{F}^{L}-\bar{\pi}_{F}^{L}-\underline{\pi}_{F}^{H}\right)$. We now make two assumptions on profits. The first simply restates Assumption 1 into this section’s notation. The second ensures that the surplus resulting from a technology sale from a low-match to a high-match firm is positive.
Assumption 3.
$\bar{\pi}_{F}^{H}>\pi_{E}>\bar{\pi}_{F}^{L}~{}\text{and}~{}\underline{\pi}_{F}^{H}>\underline{\pi}_{F}^{L}\geq 0.$
Assumption 4.
Proposition 4.
Under Assumptions 3 and 4, firm 1’s behavior in any PBE is uniquely specified. Namely, if firm 1 has a high match, it will make an acquisition and not sell the technology; if it has a low match, it will make an acquisition and sell the startup’s technology to a high-match (but not a low-match) firm 2 if and only if
$\lambda\geq\lambda_{A}(\tau)\equiv\frac{\pi_{E}-\bar{\pi}_{F}^{L}}{\underline{\pi}_{F}^{H}+\frac{\tau}{2}\left(\bar{\pi}_{F}^{H}+\underline{\pi}_{F}^{L}-\bar{\pi}_{F}^{L}-\underline{\pi}_{F}^{H}\right)}$
and do nothing otherwise.
One can verify that for $\tau=0$ the above condition reduces to the condition (1) in the main text. As $\tau$ increases, the threshold $\lambda_{A}(\tau)$ decreases, i.e., the acquisition happens for a larger set of parameters. Hence, a low-match acquirer indeed has an incentive to sell the technology to a high-match competitor because the price compensates her for any decrease in profit due to facing a more efficient competitor. Thus firms do not have an incentive to hoard technology, while the incentive to hoard talent remains. Interestingly, talent hoarding now occurs for a strictly larger set of parameters, as the option to resell technology effectively subsidizes talent hoarding.
5.2 Partial Acquisitions
This section extends the baseline model to incorporate the possibility of partial investments. That is, a firm may acquire a stake (possibly a minority stake) in the startup without integrating it. To do so, we need to specify the payoffs resulting from such partial acquisitions as well as the rights that come with partial ownership.
Formally, the ability to make partial acquisitions means that firms can, in addition to full acquihires, try to acquire a share $s\in(0,1]$ of the startup and continue to operate it as a stand-alone entity. In contrast to the baseline model, we allow for upfront and deferred payments $(p,d)$ as is standard in such transactions. [^17] In the baseline model, allowing for for upfront and deferred payments would not change anything. This allows the acquiring firm to distinguish at least partially between the investor (who only gets a part of the upfront payment) and the entrepreneur (who can also receive deferred payments). If a firm acquires a share $s$ with bid $(p,d)$, its payoff reads $\Pi_{F}+s\pi_{E}(s)-p-d$, while the entrepreneur’s payoff is $(1-s)\pi_{E}(s)+w(s)+p+d$. Here, $\pi_{E}(s)$ captures the startup’s profit net of potential wages paid to the entrepreneur as a function of the size of the external ownership. These profits accrue to the firm and the entrepreneur proportionate to their stake in the startup. Correspondingly, $w(s)$ constitutes the entrepreneur’s wage (net of effort costs) for different degrees of outside ownership. In particular, $\pi_{E}=\pi_{E}(0)+w(0)$. That is, when the entrepreneur owns the entire startup and thus obtains all of its profit and “pays herself” a net-of-effort wage, her payoff coincides with the initial payoff in the baseline model. The other firm’s payoff is unchanged at $\Pi_{F}$. When a firm acquires a stake in the startup it receives a share of profits, while the dilution of ownership gives rise to moral hazard on the entrepreneur’s side, which we capture in reduced form, only imposing the following assumption.
Assumption 5.
We assume that $\pi_{E}(s)+w(s)$ is decreasing in $s$ with $\pi_{E}(0)+w(0)>\pi_{E}(1)+w(1)$.
Assumption 5 captures the moral hazard arising when the entrepreneur no longer fully owns the startup. $\pi_{E}(s)+w(s)$ being decreasing reflects the reduced effort of the entrepreneur as a result of the agency problem. $\pi_{E}(0)+w(0)>\pi_{E}(1)+w(1)$ captures that the value of a startup fully owned by the entrepreneur is strictly higher than the value when the entrepreneur is actually an employee and has no stake in the startup.
Finally, considering control rights, we are primarily interested in the investor’s ability to block, and the entrepreneur’s ability to push through, an acquihire by the investor’s competitor. We assume that the entrepreneur can always block an acquihire, as she could sell her shares but refuse to work for the acquiring firm. If the entrepreneur would like to be acquihired and the investor does not want the acquihire to go through, the entrepreneur can (in the spirit of typical shareholder agreements) try to “drag along” the investor and force the transaction. We assume that the investor has a probabilistic chance of blocking this attempt and merely impose that the probability of successfully blocking an acquihire is increasing in the investor’s stake in the startup.
The timing of the game is as follows. In the first stage, firm 1 has the opportunity to make a bid to the entrepreneur to acquire a share $s_{1}\in(0,1]$ of the startup or make an acquihire. If the entrepreneur accepts a bid for an acquihire, the game ends. Otherwise, we move to stage 2, the ownership structure of the startup depending on whether the entrepreneur accepted firm 1’s bid. In the second stage, firm 2 can make an acquihire by making a (per-share) bid to the owner(s) of the startup, where we restrict attention to bids $(p,d)$ with $p\geq\pi_{E}(s_{1})$, so that no owner can be expropriated. If the entrepreneur does not accept the bid, the game ends. If the entrepreneur accepts, firm 1 can either also accept or try to block the transaction and succeeds in doing so with probability $\beta(s_{1})$, where $\beta$ is a weakly increasing function with $\beta(0)=0$ and $\beta(1)=1$. Nature determines whether a potential blocking attempt succeeds and the game ends.
Proposition 5.
Under Assumptions 1 and 5 and without blocking rights, i.e., $q(\beta)=0$ for all $\beta\in[0,1]$, firm 1’s behavior in any PBE is uniquely specified. Namely, if firm 1 draws a high match type, it will make an acquihire; if it draws a low type it will make an acquihire if and only if $\lambda\geq\lambda_{A}$ and do nothing otherwise.
The result in Proposition 5 shows that partial ownership of the startup is not enough to change firm 1’s behavior relative to the setting with only acquihires in Proposition 1. Indeed, since an investment is not profitable in itself and does not prevent a high-match firm 2 from making an acquihire, firm 1 will continue to either do nothing or make an acquihire, depending on the probability of a high-match firm 2 materializing. As the next result shows, it is ownership accompanied by some measure of control over the startup, which makes investments attractive to firm 1.
Proposition 6.
Under Assumptions 1 and 5 and with blocking rights, firm 1’s behavior in any PBE is uniquely specified. Namely, if firm 1 draws a high type, it will make an acquihire. If it draws a low type, there is a threshold value for $\lambda$ below which it does nothing. Above the threshold, it will do an investment and, depending on the model’s parameters, there may be an even higher threshold above which it does an acquihire.
Thus, we find that an acquihire can be more profitable than buying a startup and letting it operate independently, implying that talent hoarding persists in this extension. Further, with sufficiently strong blocking rights, an investment may constitute a viable and cheaper alternative to an acquihire. Therefore, as investments are cheaper than acquihires, the possibility of partial ownership may reduce the frequency of acquihires while increasing the frequency with which some transaction takes place. Notably, the reduction in overall profits is lower in the case of investments than in the case of low-match acquihires. Therefore, the possibility of investments is more likely to lead to inefficient market outcomes, albeit at a lower degree of inefficiency. Put differently, allowing for investments increases the extensive margin of talent hoarding but partially decreases its intensive margin.
@@ 6 Discussion and Conclusion
We have presented a simple, yet general, reduced-form model of startup acquisitions. We showed that acquihires may not only reflect firms’ desire to hire talented employees but also be rooted in an incentive to engage in inefficient talent hoarding, thus potentially warranting regulators’ attention. Further, we showed that acquihires may decrease consumer surplus and increase job volatility of acquihired employees, thereby giving further reasons for regulatory scrutiny. Finally, following the inclusion of realistic features into our baseline model such as the startup owning valuable technology that can be resold or the possibility of investing in the startup rather than acquiring it outright, talent hoarding not only persists but may even be exacerbated.
Our baseline model relies on several simplifying assumptions. As we discuss below, the main results hold even if we relax some of those assumptions. The formal analysis and results can be found in Appendix B.
@@
Uncertain Order of Moves
As noted above, talent hoarding is likely to be a problem whenever firms can identify highly impactful individuals before their competitors do so. However, when a firm becomes aware of a startup with such employees, it may not know whether its competitors are aware of the startup. This extension takes this into account by allowing for an unknown order of moves, where the order of moves is privately drawn by nature with uniform probabilities. We show that talent hoarding still emerges in that setting.
Surplus Sharing
In our baseline model, the firm has all the bargaining power and thus manages to acquire the startup at a price that leaves the entrepreneur with zero surplus. In this extension, we show that the emergence of talent hoarding does not hinge on this very asymmetric distribution of bargaining power. Instead, we show that talent hoarding emerges as long as the entrepreneur does not have too much bargaining power, as otherwise, the acquisition of the startup becomes prohibitively expensive for the firm.
@@
Appendix A Proofs
A.1 Proof of Proposition 3
@@
Absent incentives to hoard talent, only high-match firms do an acquihire. Thus, a layoff only takes place if the economy enters a downturn and the period-1 acquirer is hit by an adverse shock. Further, following a layoff, we observe the entrepreneur not being hired by the competitor only if the competitor was a low-match firm or if it was a high-match firm that got hit by an adverse shock. Finally, hiring takes place in period 1 unless both firms are low-match. Taken together, the probability of observing a layoff in period 2 is
@@
and the probability of observing the entrepreneur exiting the industry in period 2 is
@@
First, note that period-2 incentives coincide with those in period 1 absent an economic downturn, ruling out any layoffs. Following a downturn, three cases arise in period 2:
@@
To compare hiring and layoffs, we need to consider three cases.
@@
Case 1: $\lambda(1-r)(1-\gamma)>\lambda_{A}$. Then, $\lambda\geq\lambda_{A}^{1}=\lambda_{A}\frac{2}{2-\gamma\delta}$. Hence, a low-match firm will do an acquihire in both periods so no layoffs are observed, which is less than in the benchmark. Thus, irrespective of whether the economy hits a downturn, the entrepreneur is always employed without separation.
@@
Case 2: $\lambda(1-\gamma(1-r))>\lambda_{A}>\lambda(1-r)(1-\gamma)$. Hence, $\lambda>\lambda_{A}^{2}$ so firm 1 will always do an acquihire in period 1. In period 2, firm 1 will maintain employment of the entrepreneur unless it receives shock $D$. Hence, the probability of a layoff will be $\delta\gamma$, which is larger than the benchmark layoff rate $l^{*}$.
@@
Case 3: $\lambda_{A}>\lambda(1-\gamma(1-r))$. If $\lambda>\lambda_{A}^{3}$ then firm 1 will do an acquihire in period 1. Firm 1 will maintain employment in period 2 unless it was hit by a downturn and has a low match. Hence, the probability of observing a layoff is $\delta(\gamma+(1-\lambda)(1-\gamma))=\delta(1-\lambda(1-\gamma))$, which is larger than $l^{*}$. If instead $\lambda<\lambda_{A}^{3}$, then we have no talent hoarding in either stage and the equilibrium is identical to the benchmark.
@@
Finally, observe that the condition $\min\left\{\frac{\lambda_{A}}{\lambda},\frac{1-\lambda}{\lambda}\right\}>(1-r)(1-\gamma)$ implies that we are either in Case 2 or 3.
A.2 Proof of Proposition 4